OpenStudy (pucusana):

please help me,Determine the number of solutions to the following linear equation: 2x+1−3x=3−(x+2) Question 2 options: a) Zero (0) solutions b) Infinitely many solutions c) Two (2) solutions d) One (1) solution

OpenStudy (pucusana):

please help me

OpenStudy (anonymous):

hi pucusana this is the solution 2x+1-3x=3-(x+2) -x+1=3-x-2 -x+1=-x+1 so we are left with -x+1=-x+1, they basically cancel each other right. so its 0=0 now youre wondering how many solutions are there, well lets plug in some numbers EX. 1 plugin 1 2(1)+1-3(1)=3-(1+2) 2+1-3=3-1-2 0=0 EX.2 plugin 2 2(2)+1-3(2)=3-(2+2) 4+1-6=3-2-2 5-6=1-2 -1=-1 so the final is that there are infinite solutions (b)

OpenStudy (anonymous):

correct me if im wrong on something

OpenStudy (pucusana):

are you sure this is right?

OpenStudy (anonymous):

yes i just confirmed with my classmate

OpenStudy (pucusana):

nevermind i have another question.

OpenStudy (pucusana):

Determine the number of solutions for the following linear equation: 4x−8=4(x−1) Question 3 options: a) One (1) solution b) Zero (0) solutions c) Two (2) solutions d) Infinitely many solutions

OpenStudy (pucusana):

this stumped me too.

OpenStudy (anonymous):

i think vandalisk is handling this one

OpenStudy (anonymous):

I know the answer. Infinitely many solutions.(b) Because when we try to make it simplier, we miss x at all. That means that whatever x will be, we would have right equation

OpenStudy (pucusana):

oh thank you thank you!

OpenStudy (anonymous):

good job v

OpenStudy (anonymous):

just basically plugin any number into x

OpenStudy (anonymous):

any more questions?

OpenStudy (anonymous):

2x+1−3x=3−(x+2) FOIL 2(0)+1-3(0)=-3x-6 2(0)+1-3(0)=-3(0)-6 1= -6 ?